OpenStudy (anonymous):

Could someone help check my work? Calculate the scalar projection and projection of 5,3) onto (7,-2) and interpret it this projection graphically

OpenStudy (anonymous):

\[(7,-2)/ \sqrt{7^2}+(-2)^2\] =\[(7,-2) /\sqrt{53}\]

OpenStudy (anonymous):

\[(\frac{ 7 }{ \sqrt{53} }), (\frac{ -2 }{ \sqrt{53} })\]

OpenStudy (anonymous):